# thanks: Corresponding author

### v1

@@ Abstract
Due to the success of the Standard Model (SM), it is reasonable to anticipate that, the signal of new physics (NP) beyond the SM is small, and future searches for NP and precision tests of the SM will require high luminosity collider experiments. Moreover, as the precision tests of the SM advances, rarer processes with a greater number of final-state particles will require consideration, which will in turn require the analysis of a multitude of observables. As an inherent consequence of the high luminosity, the generation of a large amount of experimental data in a large feature space presents a significant challenge for data processing. In recent years, quantum machine learning has emerged as a promising approach for processing large amounts of complex data on a quantum computer. In this study, we propose a variational quantum searching neighbor (VQSN) algorithm to search for NP. As an example, we apply the VQSN in the phenomenological study of the gluon quartic gauge couplings (gQGCs) at the Large Hadron Collider. The results suggest that VQSN demonstrates superior efficiency to a classical counterpart k-nearest neighbor algorithm, even when dealing with classical data.
@@ I Introduction
In the meantime, quantum computing represents another effective approach for processing large volumes of data. Despite the current status of quantum computing as a noisy intermediate-scale quantum (NISQ) technology [37, 38, 39], there has been a notable increase in research activity within the HEP community related to this field [40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59]. Many ML algorithms can be implemented by quantum computing [60, 61, 62]. In conjunction with the advancement of quantum ML algorithms, there has been a growing interest in the deployment of quantum ML in the HEP phenomenological studies, encompassing techniques such as variational quantum classifier (VQC) [63, 64, 65, 66], quantum support vector machine (QSVM) [63, 67, 68, 69], and quantum kernel k-means (QKKM) [70], etc. In this study, we propose a variational quantum searching neighbor (VQSN) algorithm for identifying NP signals.
@@ III Variation quantum searching neighbor algorithm
For the (t-)KNN algorithm, we only incorporate the cost needed to compute the distances, so the computational complexity (cc) can be estimated as $\mathcal{O}(dNM)$ for $M$ test events. In the context of quantum computers, the number of quantum gates is typically employed as a means of assessing the cc. Even if the dataset is classical, the cc for VQSN is $n_{t}+\mathcal{O}(cN(n_{ans}+d)\sqrt{d})$, where $n_{t}$ is the cost to train the ansatz which is a constant in the view of test dataset, $n_{ans}$ is the cc of the ansatz, the cc of $U_{|\phi\rangle}^{\dagger}$ is $\mathcal{O}(d)$, and the $\mathcal{O}(\sqrt{d})$ comes from the success rate to create the state $\langle\phi|\Psi(\alpha_{i})\rangle$. With a large $M$, it can be roughly estimated that the VQSN outperforms (t-)KNN when $c(n_{ans}+d)<N\sqrt{d}$. From the principle of the algorithm, the larger $N$ and $d$, the better the classification, which means that ideally VQSN can outperform (t-)KNN even dealing with a classical dataset.
@@ IV Search of the gluon quartic gauge couplings
The training dateset has $N=65536$ events, which can be stored in $n_{q}=19$ qubits. We use the ‘hardware efficient’ ansatz [86] with the ‘shifted-circular-alternating’ entanglement layer [87], which is shown in Fig. 1. (d). The circuits are simulated using QuEST [88]. Given that $d$ is small, we do not employ the amplitude amplification, therefore the cc is about $\left((2+3l)n_{q}+2(2d-\log_{2}(d)-2)\right)cd$ for a test event. If the cc of VQSN is to be less than that of its KNN counterpart (which is $d^{\prime}N=8.5\times 10^{5}$), $l\leq 17$. In this work, we consider three cases, $l=5$, $10$, and $15$. As a comparison, we consider (t-)KNN with $K=1$ , $10$, $100$, and $1000$, with neighbors determined using Euclidean distance.
@@
This study has not addressed further improvements such as encoding the dataset using a nonlinear approach [91]. It is also possible to utilize the vast Hilbert space of quantum computers to separate the signal from the background more when mapping to the quantum feature space [92]. Furthermore, it would be beneficial to investigate the simulation of the collision process on a quantum computer. Both QSN and VQSN could be employed as a part of the simulation which are able to process data directly from a quantum computer. In conclusion, QSN and VQSN are suitable for the NP phenomenological study.

### v2

@@ Abstract
Due to the success of the Standard Model (SM), it is reasonable to anticipate that, the signal of new physics (NP) beyond the SM is small, and future searches for NP and precision tests of the SM will require high luminosity collider experiments. Moreover, as the precision tests of the SM advances, rarer processes with a greater number of final-state particles will require consideration, which will in turn require the analysis of a multitude of observables. As an inherent consequence of the high luminosity, the generation of a large amount of experimental data in a large feature space presents a significant challenge for data processing. In recent years, quantum machine learning has emerged as a promising approach for processing large amounts of complex data on a quantum computer. In this study, we propose quantum searching neighbor (QSN) and variational QSN (VQSN) algorithms to search for NP. The QSN is a classification algorithm; the VQSN, however, introduces variation to the QSN to process classical data. As an example, we apply the (V)QSN in the phenomenological study of the gluon quartic gauge couplings (gQGCs) at the Large Hadron Collider. The results suggest that VQSN demonstrates superior efficiency to a classical counterpart k-nearest neighbor algorithm, even when dealing with classical data.
@@ I Introduction
In the meantime, quantum computing represents another effective approach for processing large volumes of data. Despite the current status of quantum computing as a noisy intermediate-scale quantum (NISQ) technology [37, 38, 39], there has been a notable increase in research activity within the HEP community related to this field [40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59]. Many ML algorithms can be implemented by quantum computing [60, 61, 62]. In conjunction with the advancement of quantum ML algorithms, there has been a growing interest in the deployment of quantum ML in the HEP phenomenological studies, encompassing techniques such as variational quantum classifier (VQC) [63, 64, 65, 66], quantum support vector machine (QSVM) [63, 67, 68, 69], and quantum kernel k-means (QKKM) [70], etc. In this study, we propose a quantum searching neighbor (QSN) algorithm for identifying NP signals. In order to work in the NISQ stage, we also introduce a variational version of QSN (VQSN).
@@ III Variation quantum searching neighbor algorithm
For the (t-)KNN algorithm, we only incorporate the cost needed to compute the distances, so the computational complexity (cc) can be estimated as $\mathcal{O}(dNM)$ for $M$ test events. In the context of quantum computers, the number of quantum gates is typically employed as a means of assessing the cc. Even if the dataset is classical, the cc for VQSN is $n_{t}+\mathcal{O}(cM(n_{ans}+d)\sqrt{d})$, where $n_{t}$ is the cost to train the ansatz which is a constant in the view of test dataset, $n_{ans}$ is the cc of the ansatz, the cc of $U_{|\phi\rangle}^{\dagger}$ is $\mathcal{O}(d)$, and the $\mathcal{O}(\sqrt{d})$ comes from the success rate to create the state $\langle\phi|\Psi(\alpha_{i})\rangle$. Note that, $n_{t}$ depends on $N$ and $d$, but not on $M$. Unlike $M$ which grows with the luminosity, $N$ and $d$ remain effectively constant in practical applications, as a result, for practical applications it can be expected that $M\gg N$. By omitting $n_{t}$ because $M\gg N$, it can be estimated that the VQSN outperforms (t-)KNN even with a classical dataset when $c(n_{ans}+d)<N\sqrt{d}$. From the principle of the algorithm, the larger $N$ and $d$, the better the classification, as long as $M\gg N$ is kept, this means that ideally VQSN can outperform (t-)KNN even dealing with a classical dataset.
@@ IV Search of the gluon quartic gauge couplings
The training dateset has $N=65536$ events, which can be stored in $n_{q}=19$ qubits. We use the ‘hardware efficient’ ansatz [86] with the ‘shifted-circular-alternating’ entanglement layer [87], which is shown in Fig. 1. (d). The circuits are simulated using QuEST [88]. Given that $d$ is small, we do not employ the amplitude amplification, therefore the cc is about $\left((2+3l)n_{q}+2(2d-\log_{2}(d)-2)\right)cd$ for a test event. If the cc of VQSN is to be less than that of its KNN counterpart (which is $d^{\prime}N=8.5\times 10^{5}$), $l\leq 17$. In this work, we consider three cases, $l=5$, $10$, and $15$. As a comparison, we consider (t-)KNN with $K=1$ , $10$, $100$, and $1000$, with neighbors determined using Euclidean distance. It is found that the performance of (t-)KNN grows with $K$, but the enhancement becomes insignificant when $K>100$, for computational convenience, so we set $K$ up to a maximum of $1000$.
@@
This study has not addressed further improvements such as encoding the dataset using a nonlinear approach [91]. It is also possible to utilize the vast Hilbert space of quantum computers to separate the signal from the background more when mapping to the quantum feature space [92]. Both improvements can be applied to QSN and VQSN. Furthermore, it would be beneficial to investigate the simulation of the collision process on a quantum computer. Reading results directly from a quantum computer is resource intensive, QSN can process data within the quantum computer, avoiding this consumption. In particular, it is plausible that the training dataset originates from simulations on a quantum computer, while the test dataset is derived from classical data obtained from a collider. This scenario is well-suited to the QSN.
In conclusion, QSN and VQSN are suitable for the NP phenomenological study.